OpenStudy (anonymous):

Which shows this equation in standard form? 6 – 2(x – y) = 3x + 5 A. 5x – 2y = 1 B. 5x – 2y = –1 C. 5x + 2y = 1 D. 2x – 5y = 1 i think it's d

OpenStudy (anonymous):

no its A

OpenStudy (haleyelizabeth2017):

It's C

OpenStudy (anonymous):

no its not

OpenStudy (haleyelizabeth2017):

yes. it is

OpenStudy (anonymous):

no I had this same question before

OpenStudy (anonymous):

how do you know it c?

OpenStudy (haleyelizabeth2017):

Never mind. just listen to kyrea

OpenStudy (anonymous):

6-2(x-y)=3x+5 6-2x+2y=3x+5 6-5= 3x+2x -2y 5x-2y=1

OpenStudy (haleyelizabeth2017):

I miscalculated when I solved it :'( I'm sorry @kyrea

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

what about this one? Complete the standard form of the equation -4/3y=3-x _______x + ________y = 9

OpenStudy (haleyelizabeth2017):

-4y 3x?

OpenStudy (anonymous):

thats what i got what do you think @kyrea

OpenStudy (anonymous):

Yea its right

OpenStudy (anonymous):

ok how do i give medals

OpenStudy (anonymous):

click best response

OpenStudy (anonymous):

If you ever need help just tag me

OpenStudy (anonymous):

ok already did that will you help me with 1 more please :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the x-intercept of the line with the equation ? 1/3x+y=-15? ( __, 0)

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

i think let me make sure

OpenStudy (anonymous):

i got 2 but i see where i messed up thank you for your help @kyrea you are awesome at math :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your welcome and thx any time

OpenStudy (anonymous):

wait were you right @kyrea

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

yw